# Data-adaptive structural change-point detection via isolation

### v2

@@ 3.3 Probability of guaranteed isolation of the change-point at detection
We now return to the discussion of Section 2.3, that concerns the connection between the location of the largest difference and the guarantee that the detection of the change-point closest to it occurs in an interval in which it is isolated. The probability of this happening is given by
$\mathbb{P}\left(d_{s,e}\in\left[1,r_{1}+\frac{\tilde{\delta}_{1}}{2}-\frac{3\delta_{T}}{4n}\right]\cup\left[r_{N}-\frac{\tilde{\delta}_{N-1}}{2}+\frac{3\delta_{T}}{4n},T\right]\cup\left\{\bigcup_{j=2}^{N-1}B_{n;j}\right\}\right)$ (14)
where $B_{n;j}$ is defined in (8). Under the assumption of $\epsilon_{t}$ being i.i.d. random variables following the $\mathcal{N}(0,1)$ distribution, and in the case that $f_{t}$ is piecewise-constant, we note that in order to calculate the probability that the largest difference is exactly at the location of a change-point, we need to calculate the following:
@@
which is the probability that the observations at the change-point $r_{j}$, for any $j\in\{1,\ldots,N\}$, have a larger absolute difference than a pair of observations in the signal where no change occurs. Using (1), the fact that $\epsilon_{t}\sim\mathcal{N}(0,1)$, and that $f_{t^{\prime}+1}-f_{t^{\prime}}=0$ for $t^{\prime}\neq r_{j},j\in\{1,\ldots,N\}$, (16) can be written as
@@
$=\mathbb{P}\Biggl{(}|\alpha_{j}+Z_{2}|\geq|\tilde{Z}_{2}|\Biggr{)}$
where $\alpha_{j}=(f_{r_{j}+1}-f_{r_{j}})/\sigma$ and $Z_{2},\tilde{Z}_{2}\sim\mathcal{N}(0,2)$ are independent. In the above equation, $|\alpha_{j}+Z_{2}|$ follows the Folded Normal (FN) distribution with mean $\alpha_{j}$ and variance 2. Similarly, $|\tilde{Z}_{2}|$, follows the same distribution with mean 0 and variance 2, which reduces to the simpler case of the Half Normal (HN) distribution with variance 2. So, (17) can be equivalently written as $\mathbb{P}(Y_{1}\geq Y_{2})$ for $Y_{1}\sim FN(\alpha_{j},2)$ and $Y_{2}\sim HN(0,2)$. Denoting by $f_{Y_{1}}$ the probability density function of $Y_{1}$ and $F_{Y_{2}}$ the cumulative density function of $Y_{2}$, (17) can be written as
@@
where $t\textrm{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\rm{d}$t is the Gauss error function. Using [38] (equation (38) on p.10), we can calculate the above in terms of $\alpha_{j}$ as
$\mathbb{P}(Y_{1}\geq Y_{2})=\frac{1}{2}\biggl{[}1+\textrm{erf}\Bigl{(}\frac{\alpha_{j}}{2\sqrt{2}}\Bigl{)}^{2}\biggr{]},$ (18)
@@
so we have an explicit expression for (17). It is easy to see now that this probability goes to 1 as $\alpha_{j}$ gets larger. However, calculating or bounding from below the probability in (15), and subsequently in (14), are very difficult tasks due to the fact that $|X_{t}-X_{t-1}|$ and $|X_{t-1}-X_{t-2}|$ are dependent. Therefore, we provide results based on simulations.
@@
We want to note that when $N=1$, the interval in assumption (A1) becomes $[1,T]$, which implies that isolation of the change-point in the interval in which detection occurs is guaranteed for all possible values of the largest difference. Also, in a data sequence with $N\geq 2$ change-points, we expect the order that they are detected to be with decreasing magnitude of change. This means that the change-point with the smallest magnitude of change will be detected last, in an interval in which it is certainly isolated, so the second smallest magnitude of change in the data sequence is the one that needs to be ‘large enough’ to ensure that all change-points will be isolated in an interval where detection can occur. We performed some simulations to check what values are considered ‘large enough’.
In our simulations we consider a data sequence with $\sigma=1$ and two change-points at locations $r_{1}=20$ and $r_{2}=T-20$, with equal magnitudes of change, $\alpha=\alpha_{1}=\alpha_{2}$. It is easy to see that $\delta_{T}=20$ and $\tilde{\delta}_{1}=T-40$. We calculated the Monte Carlo estimates of (14), which is the probability that the location of the largest difference ensured that detection occurred in an interval where the change-point is isolated as $T$ increased, using 10.000 repetitions for each value of $T$. In other words, choosing $n=\frac{\delta_{T}}{2\lambda_{T}}$ for $\lambda_{T}=3$ we calculated the proportion of times that $d_{1,T}\in\left[1,\frac{T}{2}-4.5\right]\cup\left[\frac{T}{2}+4.5,T\right].$
In order to be consistent with assumption (A2), we have $\underline{f}_{T}=c\sqrt{\log T}$, where $c>0$ is a constant; in this case, $c=\frac{1}{2}$ was chosen for the sake of running the simulations and there is no other reason behind this specific choice. The results in Table 1 show that the probability of the largest difference occurring at a point that guarantees isolation in the interval where detection occurred increases with the length of the signal. This is expected as can be seen by (18) and using the fact that $\underline{f}_{T}=\mathcal{O}(\sqrt{\log(T)})$ if $\delta_{T}$ is constant, which is the case in this simulation study. It is worth noting that the proportions are high even for small $T$ and $\alpha$, which provides evidence that the probability of isolation is extremely high.
@@ 4.1 Computational complexity
Now, focusing more on the algorithms ID and DAIS, we need to highlight that the data-adaptive nature of DAIS leads to computational advancements. More specifically, due to the methodology of ID, when the data sequence under consideration has a small number of true change-points, the number of intervals that need to be checked will be close to the maximum possible for ID, which is $2\lceil\frac{T}{\lambda_{T}}\rceil$. In contrast, DAIS needs to check only $K^{\prime}$ intervals, which will be much smaller compared to ID. For example, when the signal has no change-points, using the same expansion parameter $\lambda_{T}$, DAIS checks almost half as many intervals as ID. This can be seen in the simulation results of signals (S11) and (S12) in Table 8 in Appendix B, which have 0 and 1 change-points, respectively, when compared with ID_th, which uses the same $\lambda_{T}$ as DAIS. When many change-points are present, using again the definition of $\delta_{T}$ and the inequality $\delta_{T}\geq 2n\lambda_{T}$, for which the explanation is in Section 2.3, we get $N\leq T/(2n\lambda_{T})<T/(2\lambda_{T})-1.$ So, $K^{\prime}$ will always be smaller than the maximum number of intervals that are checked by ID. However, it needs to be noted that the data-adaptivity comes at an additional, but small, computational cost, since, at each step of the algorithm, the location of the largest difference, in absolute value, needs to be re-calculated.
@@ 4.2 Parameter selection
For the choice of the positive threshold constant $C$, we ran an extensive simulation study using various signal structures and Gaussian noise. The best behaviour occurred for $C=1.7$ in the case of piecewise-constant signals and $C=2.1$ in the case of piecewise-linear signals. These are the constants that were used for the simulations in Section 5. [2] proves in Corollary 1 that, in the case of piecewise-constant signals, as $T\rightarrow\infty$, the threshold can be taken to be at most $\sqrt{3\log T}$, meaning that $C\leq\sqrt{3}\approx 1.73$; our choice does not violate this result.
Choice of the expanding parameter $\lambda_{T}$: As can be seen by the proofs of Theorems 1 and 2, in a given signal, detection will occur for any $\lambda_{T}\leq\delta_{T}/2n$, where $3/2\leq n\leq\delta_{T}/2\lambda_{T}$. Using $\lambda_{T}=1$ ensures that each change-point is isolated for as many intervals as possible, but this increases the computational time. In practice, we use $\lambda_{T}=3$, which is small enough to have very good accuracy in the detection of the change-points, but not too small which could magnify the computational cost.
@@ 5 Simulations
This section compares the performance of DAIS with state-of-the-art competitors. We only focus on algorithms that estimate both the number and the locations of the change-points in the given univariate signal in an offline manner. The competitors used in the simulation study can be found in Table 2, along with the relevant R library. For ID, WBS and SeedBS we report the results for both the information criterion and the thresholding (for the default values $C=1.56$ for ID and $C=\sqrt{2}$ for WBS and SeedBS) stopping rules. The notation is ID_ic, WBS_ic, SeedBS_ic and ID_th, WBS_th, SeedBS_th, respectively. In order to allow the competitors achieve their best performance, in some cases the inputs of the relevant functions as developed in R were adjusted. The competitors ID, WBS2, PELT and MARS were applied using the default values, while for DP_univar, the values proposed in the examples of the corresponding R function are used. For MOSUM, the function multiscale.localPrune was used, with the default values, which coincide with what is described in the simulation study of [16]. The maximum number of change-points to check for is required for the NOT algorithm and the number was chosen as $\max\{25,N+1\}$, where 25 is the default value. For MSCP, the value of the minimal window considered is chosen to be $\min\{\lceil\delta_{T}/2\rceil,20\}$, where 20 is the default value. The estimation of the standard deviation for CPOP is set to be done using the MAD method as described in Section 3. TF and SeedBS are employed based on the implementations found in https://github.com/hadley/l1tf and https://github.com/kovacssolt/SeedBinSeg/blob/master/SeedBS.R, respectively. Change-point detection using local_poly involves performing cross validation for every time series, for the optimal values to be chosen. The degree of the polynomial is set to 1, so that the fitted signals are piecewise-linear. Code on the simulations, as well as on how to implement DAIS, can be found in https://github.com/Sophia-Loizidou/DAIS. For the simulations we have used a modification of the DAIS algorithm described in Section 2, in order to make it more accurate. More specifically, DAIS restarts from the detected change-point instead of the start- and end-points of the interval where the detection occurred. This leads to an improvement of the accuracy of the algorithm without affecting the speed. This is also our recommendation on how to use the algorithm in practice.
@@
ID [2] IDetect
PELT [30] changepoint
WBS [23] breakfast
WBS2 [25] breakfast
Piecewise-constant NOT [6] not
MOSUM [37] mosum
MSCP [33] mscp
@@
DP_univar [49] changepoints
ID [2] IDetect
CPOP [21] cpop
Continuous, NOT [6] not
piecewise-linear MARS [22] earth
@@
TS [35] trendsegmentR
local_poly [54] changepoints
@@
where $n_{s}=\max_{j=1,\ldots,N}\tilde{\delta}_{j}$ is the length of the largest segment between successive change-points. The Hausdorff distance, $d_{H}$, is given for all signals except for signals that do not contain any change-points; in such cases, $d_{H}$ is uninformative. Finally, the average computational time for each method is also reported. The top performing algorithm, in terms of accuracy in the number of change-points detected, is presented in bold, along with all methods that performed within $5\%$. The signals used can be found in Appendix A and some further simulation results are in Appendix B.
The signals included in this section contain more than one change-point. For signals with no or one change-point, please refer to Tables 8 and 10 in Appendix B. The signal (S1) (Table 3) is a particularly difficult structure for most algorithms because it includes a pair of consecutive change-points that shift the data sequence to opposite directions, which masks the existence of the change-points. Detection is more difficult when the change-points are close to each other and their isolation is unlikely or occurs only in a limited number of intervals, which is the case in this signal. The best performing method for (S1) is DAIS, having at least a $7.5\%$ improvement in the number of times the correct number of change-points is detected compared to all the other algorithms. The reason why DAIS performs very well is that it will most likely start the expansion around the change-points, as is discussed in Section 3.3, and so it has an advantage in detecting them, as the power of the contrast function $C_{s,e}^{b}(\boldsymbol{X})$ is maximized when $b$ is at the midpoint of $[s,e]$. For signals (S2) and (S3), the majority of the algorithms perform well; however, WBS_th and SeedBS_th tend to overestimate and MSCP seems to underestimate the number of the change-points, while DP_univar underestimates in signal (S2) and overestimates in signal (S3). For signals (S4), (S5) and (S6) in Tables 3 and 4, some methods struggle to detect the change-points, which are really close to each other. The signal (S4) includes change-points of different magnitudes and with different distances between them. ID_th and ID_ic perform well in different scenarios, but neither is consistent, which is expected as the thresholding approach is more appropriate for a large number of change-points, while the information criterion is better suited for a smaller number. Even though MOSUM performs well for signals with a small number of change-points, it fails to detect the change-points for (S4), (S5) and (S6). WBS2 performs well but is more computationally expensive compared to DAIS, with DAIS being 10-100 times faster than WBS2 in the simulations. Based on the simulations, we conclude that overall, DAIS performs very well in various different piecewise-constant signal scenarios, involving the number of change-points, the distance between them, as well as the magnitudes of the changes, when most competitors struggle to accurately detect the changes in some of the cases tested.

### v3

@@ 3.3 Probability of guaranteed isolation of the change-point at detection
In this section, we work under Assumptions (A2) and (A3), for piecewise-constant and continuous, piecewise-linear signals, respectively. We return to the discussion of Section 2.3, that concerns the connection between the location of the largest difference and the guarantee that the detection of the change-point closest to it occurs in an interval in which it is isolated. The probability of this happening is given by
$\mathbb{P}\left(d_{s,e}\in\left[1,r_{1}+\frac{\tilde{\delta}_{1}}{2}-\frac{3\delta_{T}}{4n}\right]\cup\left[r_{N}-\frac{\tilde{\delta}_{N-1}}{2}+\frac{3\delta_{T}}{4n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{s,e}^{\max}}\right]\cup\left\{\bigcup_{j=2}^{N-1}B_{n;j}\right\}\right)$ (19)
where $B_{n;j}$ is defined in (10) and the union over $j$ is the empty set for $N\leq 2$. Under the assumption of $\epsilon_{t}$ being i.i.d. random variables following the $\mathcal{N}(0,1)$ distribution, and in the case that $f_{t}$ is piecewise-constant, we note that in order to calculate the probability that the largest difference is exactly at the location of a change-point, we need to calculate the following:
@@
which is the probability that the observations around the change-point $r_{J}$ have a larger absolute difference than a pair of observations in the signal where no change occurs. The focus is put on $t\notin\{r_{J}-1,r_{J},r_{J}+1\}$ such that $|X_{r_{J}+1}-X_{r_{J}}|$ and $|X_{t+1}-X_{t}|$ are independent. Using (1), the fact that $\epsilon_{t}\sim\mathcal{N}(0,1)$, and that $f_{t^{\prime}+1}-f_{t^{\prime}}=0$ for $t^{\prime}\neq r_{J}$, (21) can be written as
@@
$=\mathbb{P}\Biggl{(}\left|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{f_{r_{J}+1}-f_{r_{J}}}{\sigma}}+Z_{2}\right|\geq|\tilde{Z}_{2}|\Biggr{)}$
where $Z_{2},\tilde{Z}_{2}\sim\mathcal{N}(0,2)$ are independent. In the above equation, $|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(f_{r_{J}+1}-f_{r_{J}}\right)/\sigma}+Z_{2}|$ follows the Folded Normal (FN) distribution with mean ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(f_{r_{J}+1}-f_{r_{J}}\right)/\sigma}$ and variance 2. Similarly, $|\tilde{Z}_{2}|$, follows the same distribution with mean 0 and variance 2, which reduces to the simpler case of the Half Normal (HN) distribution with variance 2. So, (22) can be equivalently written as $\mathbb{P}(Y_{1}\geq Y_{2})$ for $Y_{1}\sim FN({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(f_{r_{J}+1}-f_{r_{J}}\right)/\sigma},2)$ and $Y_{2}\sim HN(0,2)$. Denoting by $f_{Y_{1}}$ the probability density function of $Y_{1}$ and $F_{Y_{2}}$ the cumulative distribution function of $Y_{2}$, (22) can be written as
@@
where $t\textrm{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\rm{d}$t is the Gauss error function. Using [47] (equation (38) on p.10), we can calculate the above as
$\mathbb{P}(Y_{1}\geq Y_{2})=\frac{1}{2}\biggl{[}1+\textrm{erf}\left(\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{r_{J}+1}-f_{r_{J}}}}{2\sqrt{2}\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma}}\right)^{2}\biggr{]},$ (23)
@@
so we have an explicit expression for (22). It is easy to see now that this probability goes to 1 as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(f_{r_{J}+1}-f_{r_{J}}\right)}$ gets larger. However, calculating or bounding from below the probability in (20), and subsequently in (19), are not as straightforward to handle tasks due to the fact that $|X_{t}-X_{t-1}|$ and $|X_{t-1}-X_{t-2}|$ are dependent. We provide results based on simulations.
@@
We want to note that when $N=1$, the interval in Assumption (A1) becomes $[1,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{s,e}^{\max}}]$, which implies that isolation of the change-point in the interval in which detection occurs is guaranteed for all possible values of the location of the largest difference. Also, in a data sequence with $N\geq 2$ change-points, we expect the order that they are detected to be with decreasing magnitude of change. This means that the change-point with the smallest magnitude of change will be detected last, in an interval in which it is certainly isolated, so the second smallest magnitude of change in the data sequence is the one that needs to be ‘large enough’ to ensure that all change-points will be isolated in an interval where detection can occur. For this reason, the following simulation study considers data sequences with two change-points.
Before investigating (19) further, it is worth emphasizing that for any location of the largest difference, isolation of the change-point is guaranteed for at least one interval if $\lambda_{T}=1$. However, as explained in Section 2.3, in order to guarantee that detection will occur, the endpoints of the interval have to be at a distance of at least $\frac{\delta_{T}}{2n}$ from the change-point, and Assumptions (A2) or (A3) need to be satisfied.
For the simulation study, we focus on piecewise-constant signals. More specifically, we consider a data sequence with $\sigma=1$ and two change-points at locations $r_{1}=10+\lfloor\log T\rfloor$ and $r_{2}=T-10-\lfloor\log T\rfloor$, with equal magnitudes of change, which are chosen to be equal to $\sigma$. These values ensure that Assumption (A2) is satisfied and are also aligned with the conclusion of the consistency theorem, Theorem 1, which states that $\delta_{T}$ has to be at least of order $O(\log T)$ in order to be able to match the estimated change-points with the true ones. We use $n=3/2$, which is the most unfavorable choice, as it minimizes the total length of the union of the intervals in (19), which becomes $\left[1,\frac{T}{2}-\delta_{T}/2\right]\cup\left[\frac{T}{2}+\delta_{T}/2,T-1\right]$. In Table 1, the proportion of times that $d_{1,T}$ lies in this union of intervals is reported as $T$ increases, using 10,000 repetitions for each value of $T$. The results show that the probability of the largest difference occurring at a point that guarantees isolation in the interval where the change-point will be detected increases with the length of the signal. This is expected in this example, as the length of the interval in which $d_{1,T}$ should not lie is equal to $\delta_{T}$, and therefore, its proportion to the length, $T-1$, of the whole differenced sequence goes to 0.
@@ 4.1 Computational complexity
Now, focusing more on the algorithms ID and DAIS, we need to highlight that the data-adaptive nature of DAIS leads to computational advancements. More specifically, due to the methodology of ID, when the data sequence under consideration has a small number of true change-points, the number of intervals that need to be checked will be close to the maximum possible for ID, which is $2\lceil\frac{T}{\lambda_{T}}\rceil$. In contrast, DAIS needs to check only $K^{\prime}$ intervals, which will be much smaller compared to ID. For example, when the signal has no change-points, using the same expansion parameter $\lambda_{T}$, DAIS checks almost half as many intervals as ID, which leads to substantial computational gain. This can be seen in the simulation results of Signals (S12) and (S13) in Table S1, which have 0 and 1 change-points, respectively, when compared with ID_th, which uses the same $\lambda_{T}$ as DAIS. When many change-points are present, using again the definition of $\delta_{T}$ and the inequality $\delta_{T}\geq 2n\lambda_{T}$, for which the explanation is in Section 2.3, we get $N\leq T/(2n\lambda_{T})<T/(2\lambda_{T})-1.$ So, $K^{\prime}$ will always be smaller than the maximum number of intervals that are checked by ID. However, it needs to be noted that the data-adaptivity comes at an additional, but small, computational cost, since, at each step of the algorithm, the location of the largest difference, in absolute value, needs to be re-calculated.
@@ 4.2 Parameter selection
For the choice of the positive threshold constant $C$, we ran an extensive simulation study using various signal structures and Gaussian noise. The best behavior occurred for $C=1.7$ in the case of piecewise-constant signals and $C=2.1$ in the case of piecewise-linear signals. These are the constants that were used for the simulations in Section 5. However, our algorithm performs well for a range of values for $C$. Simulation results for $C\in\{1.5,1.6,1.7,1.8,1.9\}$ in the case of piecewise-constant and $C\in\{1.9,2,2.1,2.2,2.3\}$ in the case of continuous, piecewise-linear signals are provided in the supplementary material in S2 in Tables S6 and S7. [3] proves in Corollary 1 that, in the case of piecewise-constant signals, as $T\rightarrow\infty$, the threshold can be taken to be at most $\sqrt{3\log T}$; our choice of 1.7 for the threshold constant, $C$ as in (24), does not violate this result.
Choice of the expanding parameter $\lambda_{T}$: As can be seen by the proofs of Theorems 1 and 2, in a given signal, detection will occur for any $\lambda_{T}\leq\delta_{T}/2n$, where $n\geq 3/2$ is a constant. Using $\lambda_{T}=1$ ensures that each change-point is isolated for as many intervals as possible, but this increases the computational time. In practice, we use $\lambda_{T}=3$, which is small enough to have very good accuracy in the detection of the change-points, but not too small which could magnify the computational cost. Our algorithm is quite robust to changes in the value of $\lambda_{T}$; for a small simulation study in the case that $\lambda_{T}\in\{1,3,5,10,15,20\}$, see Table S5. If the value of $\lambda_{T}$ is large, then there is a risk of underestimating the number of change-points in cases where the distance between change-points is smaller than $\lambda_{T}$.
@@ 5 Simulations
This section compares the performance of DAIS with state-of-the-art competitors. We only focus on algorithms that estimate both the number and the locations of the change-points in the given univariate signal in an offline manner. The competitors used in the simulation study can be found in Table 2, along with the relevant R packages. For ID, WBS and SeedBS we report the results for both the Information Criterion (IC) and the thresholding stopping rules. The notation is ID_ic, WBS_ic, SeedBS_ic and ID_th, WBS_th, SeedBS_th, respectively. Regarding the threshold constant, the default value $C=1.15\sqrt{2}$, as in the breakfast R package ([1]), was taken in the piecewise-constant case, while $C=1.4\sqrt{2}$ was used for ID in the continuous, piecewise-linear case). We highlight, that even though for SeedBS we used code (that is partly based on the wbs R package ([8])) from the relevant GitHub repository, we decided, for a fair comparison, to keep the well-calibrated threshold constant of the breakfast R package. This choice led to a substantially better performance for SeedBS compared to results that were relied on the constant used in the relevant GitHub code. Furthermore, for the NOT method of [7] results are presented based on the Schwarz information criterion, while for the WBS2 method of [28] results are provided when the algorithm is combined with the Steepest Drop to Low Levels (SDLL) model selection procedure, introduced also in [28]. In order to allow the competitors achieve their best performance, in some cases the inputs of the relevant functions, as developed in R, were adjusted. The competitors ID, WBS, WBS2, PELT and MARS were applied using the default values based on the relevant R packages provided in Table 2. For DP_univar, the values proposed in the examples of the corresponding R function are used. For MOSUM, the function multiscale.localPrune was used, with the default values for all the parameters apart from the minimal mutual distance of change points that can be expected (the d.min input argument) which was set equal to 3 instead of 10. This enhances MOSUM’s performance compared to the default d.min = 10 value because in our simulation study, we use signals with $\delta_{T}<10$. The maximum number of features allowed, $K_{\rm max}$, is required for the function that employs the Information Criterion model selection procedure within the breakfast R package. Therefore, for the NOT, WBS_ic and ID_ic methods we needed to choose $K_{\rm max}$. The default value in the relevant R function is $K_{\rm max}=\min\{100,T/\log(T)\}$. In signals, where the true number of change-points surpasses this default value, we set $K_{\rm max}=\lceil T/\delta_{T}\rceil$, where $\delta_{T}$ is the minimum distance between consecutive change-points in the signal tested. For MSCP, the value of the minimal window considered is chosen to be $\min\{\lceil\delta_{T}/2\rceil,20\}$, where 20 is the default value. SMUCE requires Monte Carlo simulations for calculating $q$, the vector of critical values at level $\alpha$, which is the asymptotic probability of overestimating the number of change points. The time for these simulations, which can be quite high, is not included in the values reported below and we use the default choice, which is $\alpha=0.5$. [41] proposes an implementation for MsFPOP and MsPELT, which are extensions of FPOP and PELT to the multiscale penalty proposed by [58]. As the results for the number and the locations of the estimated change-points are the same (they only differ at their computational times), we only report the results of MsFPOP. The calibration constants are set to the values chosen by simulations in [41] such that the percentage of the false positives is less than 0.05. The estimation of the standard deviation for CPOP is set to be done using the MAD method as described in Section 3. TF, SeedBS and MsFPOP are employed based on the implementations found in https://github.com/hadley/l1tf, https://github.com/kovacssolt/SeedBinSeg/blob/master/SeedBS.R and https://github.com/aLiehrmann/MsFPOP, respectively. Change-point detection using local_poly involves performing cross validation for every time series, for the optimal values to be chosen. The degree of the polynomial is set to 1, so that the fitted signals are piecewise-linear. Code on the simulations, as well as on how to implement DAIS, can be found in https://github.com/Sophia-Loizidou/DAIS. For the simulations we have used a modification of the DAIS algorithm described in Section 2, in order to make it more accurate. More specifically, DAIS restarts from the detected change-point instead of the start- and end-points of the interval where the detection occurred. This leads to an improvement in the accuracy of the algorithm without affecting the speed. This is also our recommendation on how to use the algorithm in practice.
@@
ID [3] breakfast ([1])
PELT [35] changepoint ([34])
WBS [26] breakfast ([1])
WBS2 [28] breakfast ([1])
NOT [7] breakfast ([1])
Piecewise-constant MOSUM [21] mosum ([45])
MSCP [39] mscp ([38])
@@
DP_univar [60] changepoints ([62])
SMUCE [24] stepR ([49])
MsFPOP [41] -
ID [3] breakfast ([1])
CPOP [23] cpop ([22])
Continuous, NOT [7] breakfast ([1])
piecewise-linear MARS [25] earth ([46])
@@
TS [42] trendsegmentR ([43])
local_poly [66] changepoints ([62])
@@
where $n_{s}=\max_{j=1,\ldots,N}\tilde{\delta}_{j}$ is the length of the largest segment between successive change-points. The Hausdorff distance, $d_{H}$, is given for all signals except for signals that do not contain any change-points; in such cases, $d_{H}$ is uninformative. Finally, the average computational time for each method is also reported. The top performing algorithm, in terms of accuracy in the number of change-points detected, is presented in bold, along with all methods that performed within $5\%$. The lowest MSE and $d_{H}$ values are also in bold, as well as all values that are within 10% of the lowest ones. The signals used can be found in Appendix A and further simulation results are in the supplementary material.
The signals included in this section contain more than one change-point. For signals with no or one change-point, please refer to Tables S1 and S3 in the supplementary material. Due to their difficult structure, we first put the focus of the discussion on Signals (S1) and (S2).
Signal (S1) (Table 3) is a particularly difficult structure for most algorithms because it includes a pair of consecutive change-points that shift the data sequence to opposite directions, which masks the existence of the change-points. Signals with this type of difficult settings make the appreciation of the isolation aspect more apparent; for a relevant discussion of the importance of localization in such signals see also Section 2 in [26]. Therefore, detection is harder when the change-points are close to each other, ‘work against each other’, and their isolation, while of great importance, is for some algorithms unlikely or, for others, occurs only in a limited number of intervals. The best performing methods in (S1), in terms of accuracy for the estimated number of change-points, are DAIS and ID_th; this is unsurprising based on the fact that these two methods are expected to certainly achieve isolation of the change-points prior to their detection.
Signal (S2) contains two change-points at a distance of just 5, at locations $r_{1}=30$ and $r_{2}=35$. The magnitudes of the changes regarding $r_{1}$ and $r_{2}$ are 2.3 and 5.7, respectively. Due to the large magnitude of change for $r_{2}$, DAIS starts checking for a change-point around $r_{2}$, which can be detected while it is still isolated in a small interval (due to the distance with $r_{1}$ being small). Then, $r_{1}$, which has a smaller magnitude of change, is also detected while isolated. In contrast to this, algorithms that are also based on thresholding and check the data sequence in a sequential way (while attempting to obtain change-point isolation), either through keeping one of the end-points fixed (for example, ID) or through a moving sum approach (for example, MOSUM) may struggle with such a signal. This is because, in such algorithms, $r_{1}$ will appear first while scanning the data, and if not detected while still isolated, then there is the risk of $r_{2}$ being first detected (due to its larger magnitude of change) before $r_{1}$, in an interval that also contains $r_{1}$. Discarding the data points already checked (as, for example, in ID), or carrying on the moving sum process further away from $r_{1}$ (as, for example, in MOSUM), will then mean that $r_{1}$ can not get detected in any following steps of such algorithms. Moving now away from thresholding based approaches, we highlight that WBS2 (which is based on the SDLL model selection procedure), as well as many optimization based algorithms, more specifically, ID_ic, WBS_ic, NOT, SeedBS_ic, and MsFPOP, have a very good behavior in (S2).
Regarding the results for Signals (S3) and (S4) as shown in Table 4, many algorithms, including our proposed method DAIS, perform very well. However, ID_ic, WBS, SeedBS, and NOT tend to slightly overestimate the true number of change-points. In contrast, the MSCP algorithm seems to underestimate the number of the change-points in those two signals. DP_univar and MsFPOP exhibit heavy underestimation in Signal (S3), while they heavily overestimate the true number of change-points in (S4). SMUCE underestimates in Signal (S3) and performs very well in (S4). For Signals (S5), (S6) and (S7) shown in Tables 4 and 5, a great number of methods struggle to detect the change-points, due to the limited spacings between them. Signal (S5) includes change-points of different magnitudes and with slightly different distances between them. DAIS, ID, WBS_th, WBS2, PELT, and MsFPOP exhibit very good performance in the signal both in terms of accuracy (with respect to the number and the locations of the estimated change-points) and speed. Signals (S6) and (S7) consist of a large number of regularly occurring change-points. DAIS, ID_th, WBS2, SeedBS_th, and MsFPOP exhibit a very good behaviour in such complex structures. Notice that algorithms which were tested with both the thresholding and IC stopping rules, such as ID, WBS, and SeedBS, tend to perform well in general but neither is consistently accurate in terms of the choice of the stopping rule in the various scenarios tested that involve large deviations on the number of change-points as well as their distance. For example, notice the dependency on the model selection method (either IC or thresholding) when the performances related to ID, SeedBS, and WBS for IC and thresholding in signal (S2), are compared to their performances in Signals (S6) and (S7). Such model selection based behaviors are expected as the thresholding approach is more appropriate for a large number of change-points, potentially close to each other, while the information criterion is better suited for a moderate number of change-points with larger spacings. This difference in accuracy between the threshold- and SIC-based stopping rules is what motivated in [3] the introduction of a hybrid version of ID that combines these two stopping rules. Regarding the MOSUM algorithm, it performs well for signals with a small number of change-points; however it struggles in Signals (S5), (S6) and (S7), which are more complex structures involving a large number of change-points with limited spacings. WBS2 is a very well performing algorithm in terms of accuracy, but is more computationally expensive compared to DAIS, with DAIS being at least one order of magnitude faster than WBS2 in the simulations. Based on the simulations, we conclude that overall, DAIS performs very well in various different piecewise-constant signal scenarios regarding the number of change-points, the distance between them, as well as the magnitudes of the changes, when most competitors struggle to accurately detect the changes in some of the cases tested.